Lenows Drug Stores and Halls Pharmaceuticals are competitors in the discount drug chain store business. The separate capital structures for Lenow and Hall are presented here Lenow Hall Debt @ 10% Common stock, $10 par Total Common shares $200,000 100,000 $300,000 10,000 $100,000 Debt@ 10% 200,000 Common stock, $10 par S300,000 Total 20,000 Common shares a. Complete the following table given earnings before interest and taxes of $20,000, $30,000, and $120,000. Assume the tax rate is 30 percent. (Round your answers to 2 decimal places.) What is the relationship between the EPS of the two firms? EBIT | $ 20,000 $300,000 $30,000 $300,000 $ 120,000 $300,000 Total assets EBIT/TA | % | Lenow EPS Hall EPSb-1. What is the EBIT/TA rate when the firms have equal EPS? EBIT/TA rate b-2. What is the cost of debt? Cost of debt b-3. State the relationship between earnings per share and the level of EBIT. EPS is unaffected by financial leverage when the pre-tax return on assets (EBIT/TA) the cost of debt c. If the cost of debt went up to 12 percent and all other factors remained equal, what would be the break-even level for EBIT? Break-even level
Answer #1

Please provide rating...

Answer a.
i ii iii=i/ii
EBIT Total asset EBIT/TA Lenow EPS Hall EPS relationship
20000 300000 6.67% 0.35 0 Pretax return is lower than cost of debt therefore Hall EPS is lower compared to Lenow
30000 300000 10.00% 0.7 0.7 Pretax return is equal to the cost of debt therefore both EPS are equal
120000 300000 40.00% 3.85 7 pretax return is higher than cost of debt therefore Lenow EPS is lower than Hall EPS
Computation of EPS
Lenow Hall
i EBIT 20000 30000 120000 20000 30000 120000
ii interest 10000 10000 10000 20000 20000 20000
iii=i-ii EBT 10000 20000 110000 0 10000 100000
iv=iii*.7 PAT=EBT*(1-0.3) 7000 14000 77000 0 7000 70000
v Share out standing 20000 20000 20000 10000 10000 10000
vi=iv/v EPS 0.35 0.7 3.85 0 0.7 7
Answer b.-1
at 10% EBIT/TA both the firm has equal EPS
Answer b.-2
Cost of debt =10%
Answer b.-3
EPS is unaffected by financial leverage when pre-tax return on asset (EBIT/TA) is equal to the cost of debt.
Answer c
Break even level if debt cost increased to 12%
At break even (with 12% cost of debt) we should have EBIT/TA = 12%
EBIT = 12%* TA
EBIT = 12%*300000
EBIT = 36000
Break even EBIT = 36000